OpenStudy (anonymous):

differentiate and simplify: y=(3x)/(x^2 + 1)

zepdrix (zepdrix):

\[\Large y=\frac{3x}{x^2+1}\]Quotient Rule!\[\Large y'=\frac{\color{royalblue}{(3x)'}(x^2+1)-(3x)\color{royalblue}{(x^2+1)'}}{(x^2+1)^2}\]

zepdrix (zepdrix):

So there is our quotient rule setup. We'll need to take the derivative of the blue parts. Any confusion? :o

OpenStudy (anonymous):

not yet

OpenStudy (anonymous):

y'= (3) * (x^2+1) - 2x * 3x / (x^2 +1)^2

zepdrix (zepdrix):

mmmmm looks good.